The fog collector doors were open and there was a sample but it was too small to collect. The sample was under 5ml in volume. The collector was cleaned in place and the sample jar was rinsed with DI water and placed back onto the collector to take another sample.
At 10:32 am we took a 200ml sample and a 15 ml sub-sample.
Dismantled, cleaned collector, and took a field blank.
Then the collector was restarted. Sample as well as blank were all treated and refrigerated within the hour.
Today we collected a 50ml fog sample. The sample was clean with no bugs or dirt in it, although there were quite a number of moths stuck in the strings of the collector. The sample was preserved with HCL but was not split as it was not greater than 50ml. The sample was then refrigerated. The collector was taken apart and cleaned thoroughly. the collector was then put back together and a 200ml blank was taken. The blank was preserved with HCL and then promptly refrigerated. The system was shut off because our lab is closed on Sundays. We will resume data collection on Monday.
We collected a 140ml fog sample today! There were some bugs in the sample but other than that it was a good sample. We then conducted a sub sample for anion analysis. The fog collector was taken apart, washed, then placed in an acid bath for one hour. Afterwards the collector was put back together and a new sample jar was placed.